OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 10x + 25 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (lgbasallote):

solve for the discriminant \[b^2 - 4ac\] \[(10)^2 - 4(1)(25)\] \[100 - 4(25)\] \[100 - 100\] \[0\]

OpenStudy (lgbasallote):

so how many solutions?

OpenStudy (anonymous):

SO ITS NO REAL SOLUTIONS

OpenStudy (lgbasallote):

b^2 - 4ac > 0 two real solutions b^2 - 4ac = 0 one real solution b^2 - 4ac < 0 no real solution

OpenStudy (lgbasallote):

so what do you think?

OpenStudy (anonymous):

x² + 10x + 25 = 0 We have this : b² - 4ac = 10² - 4.25 = 0 => B.Two identical rational solutions

OpenStudy (anonymous):

so its what ashley said! lol

OpenStudy (anonymous):

Lol yup!

OpenStudy (anonymous):

yes its B

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thank you everyone.. you guys and girls should help me on my other one im about to post lol

OpenStudy (anonymous):

Your welcome. & why nott (: